# Tensor networks for quantum computing

### v1

@@ Abstract
In the rapidly evolving field of quantum computing, tensor networks serve as an important tool due to their multifaceted utility. In this paper, we review the diverse applications of tensor networks and show that they are an important instrument for quantum computing. Specifically, we summarize the application of tensor networks in various domains of quantum computing, including simulation of quantum computation, quantum circuit synthesis, quantum error correction, and quantum machine learning. Finally, we provide an outlook on the opportunities and the challenges of the tensor-network techniques.
@@ I Introduction
Simulation of quantum computation (Section III) Unstructured TNs, matrix product states (MPS), projected entangled pair states (PEPS), multi-scale entanglement renormalization ansatz (MERA), density matrix renormalization group (DMRG), various renormalization methods Simulation of quantum circuits, analog quantum processors, boson sampling, quantum algorithm benchmarking, quantum many-body physics simulation Simulation complexity reduction, efficiency in representation of some quantum states Representing highly-entangled quantum states, simulating long-time dynamics, simulating deep quantum circuits
@@
Quantum error correction (QEC) (Section V) Tensor-network decoders, TN inspired quantum error-correcting codes Tensor-network decoders of error-correcting codes Efficiency and interpretability in decoding, interpretability in creating and studying error-correcting codes Computational resources overhead, speed, lack of open-source software
Quantum machine learning (QML) (Section VI) Variational tensor-network circuits as quantum neural networks (QNNs) Generative and discriminative models, quantum data classification, quantum data encoding Analytical interpretability, qubit efficiency, noise resilance, absence of barren plateaus, potential reduction in parameters number, avoiding costly TN contraction on classical computer Requirement for nonlocal interactions in quantum hardware for many TN circuits, lack of theoretical guarantees for advantage
@@
Tensor networks (TNs) have become a useful tool in many areas of physical and mathematical sciences, especially in the field of quantum information science. The interest in quantum computing (QC) [1, 2] has driven a lot of the development in TNs because they are used to represent and manipulate quantum states and processes.
@@
Additionally, TNs provide a general framework of analyzing mathematical objects frequently encountered in quantum information science, which makes them attractive tools to address many other challenges faced by QC. Table 1 provides an overview of TN applications in QC that we will discuss in this paper and spans four subdomains: quantum circuit simulations, quantum circuit synthesis (QCS), quantum error correction (QEC), and quantum machine learning (QML). The table outlines the specific methods used, their primary applications, key advantages, and the challenges they encounter.
We structure the rest of the paper as follows. Section 2 introduces tensor networks, key building blocks, and commonly used methods. Section 3, 4, 5, and 6 discuss the aforementioned domains of application. Finally, we conclude the paper with discussions on the overall advantages of tensor networks, and provide an outlook of the future on how tensor networks may benefit quantum computing.
@@ II Tensor-network methods
where $\Sigma$ is a diagonal matrix of the singular values and could be absorbed into $U$ or $V$, and $U,V$ are isometries ($U^{\dagger}U=I,V^{\dagger}V=I$). This decomposition presents an opportunity for approximate representation of the original tensor by trimming the singular values. Further, the Eckart–Young theorem [11] states that for approximations with a fixed rank, the solution provided by SVD is optimal. There exist other matrix decompositions which we do not discuss for the sake of brevity [12].
Graphically, the SVD decomposition is shown in Figure 1(e). There, isometries are represented as orangle nodes with directional bonds shown in Figure 1(f). Unitaries, being a special case of isometries, will be represented as magenta nodes with directional bonds as shown in Figure 2.
II.1 Architectures
Figure 2: Different types of TNs: (a) matrix product state, (b) multiscale entanglement renormalization ansatz with the periodic boundaries. The orange circles represent isometric tensors and the magenta circles represent unitary disentanglers. (c, d) projected entangled pair states in 2D and 3D.
The matrix product state (MPS) is the simplest yet the most widely used architecture. In general, the amplitudes of an $N$-body quantum wavefunction form an $N$th-order tensor, and the MPS provides an approximate representation of this tensor as
@@ II.1 Architectures
which is a contraction of the $\Gamma_{i}$ tensors, each corresponding to a site in the system as shown in Fig. 2 (a). Here, $\chi$ is the parameter called the bond dimension which controls the accuracy of compression. This TN is especially effective for simulating one-dimensional (1D) quantum systems. It follows the area law of entanglement [13], but has been used for other systems as well [14, 9]. Its contraction cost is polynomial in the system size, meaning the physical observables can be evaluated efficiently. This concept has independently emerged in computational mathematics, where it is called the tensor train [15].
Projected entangled pair states (PEPS), shown in Fig. 2 (c) and Fig. 2 (d), aim at generalizing the MPS to higher dimensions, making them capable of reproducing large entanglement and suitable for simulating high dimensional quantum systems. However, algorithms for PEPS are typically much more computationally demanding [5, 16].
The multiscale entanglement renormalizaton ansatz (MERA), shown in Fig. 2 (b), is designed to capture the entanglement structure of quantum systems across different length scales and can reproduce logarithmic violation of the area law.
@@ II.2 Manipulation of TNs
While TNs can be used to represent a broad range of mathematical objects in quantum information science, their popularity stems primarily from their power as a tool for obtaining meaningful, interpretable quantities such as operator expectation values, reduced density matrices, amplitudes, samples, and more. Due to the high interest in the TN techniques, a plethora of software libraries suitable for different types of TN manipulations are available, such as cuQuantum [19], Cotengra [20], iTensor [21], and many more [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@ II.2.1 Update
In addition to the DMRG algorithm, gradient-based optimization methods can also be employed if the gradient of the objective function is accessible, for instance, through automatic differentiation [38, 39]. Additionally, Riemannian optimization [40, 41] can be particularly helpful when one needs to maintain the isometric properties of the TN [42, 43].
@@
One may also consider the time evolution of a quantum state or an operator in the TN representation. For discrete time steps, this can be done by applying local gates or Kraus operators through contraction, and restore the original form of the TN with SVD. These approaches are called time-evolving block decimation (TEBD). Continuous time evolution under a Hamiltonian can be approximated by discrete Trotter steps and the discrete method can be applied if the Hamiltonian is sufficiently short ranged. For long-range Hamiltonians, TEBD becomes inefficient, and methods such as MPO $W^{II}$, TDVP, and Krylov approaches are preferred; see [46] for a recent review. Krylov methods use Lanczos diagonalization to compute the action of the time evolution operator, while TDVP evolves the state directly on the MPS manifold for fixed bond dimension [47, 48, 49, 50].
@@ II.2.2 Contraction
Finding the optimal contraction scheme is itself an NP-complete problem [52, 53]. However, much progress has recently been made using heuristic approaches targeting the total cost of the contraction operation. Specifically, these methods have been based on recursive graph partitioning [54, 20, 55], simulated annealing [56, 57] and now reinforcement learning [58, 59]. The extraordinary sensitivity of the cost to the contraction tree quality has led to improvements by many orders of magnitude for some problems such as, for example, quantum circuit simulation.
For exact simulation of quantum circuits, two additional techniques are central to achieving a state-of-the-art performance. First, whilst the time cost of a contraction might be acceptable, the space cost in terms of intermediate memory might still be enormous. Slicing (also known as cutting or projecting) splits a contraction into many smaller independent contractions [60, 61, 62, 63, 64], each of which for example can fit on a GPU and be performed in a massively parallel way [65, 56]. Depending on the geometry, this can sometimes be performed with very little overhead. Secondly, if one wants to evaluate many related tensor networks differing only in some entries (for example, a set of amplitudes), one can employ multi-contraction [66, 7] to cleverly avoid repeating the same computation. The largest exact simulations of quantum circuits use both of these techniques in tandem [7, 67, 68, 69].
@@ III.1 Gate-based quantum computation
A good example is the simulation of quantum circuits used to simulate many-body physics, such as the experiment on the dynamics of a kicked Ising model [87]. While the problem was considered initially intractable using state-vector MPS and isometric-TN [88] approaches, it was later demonstrated that TN belief propagation [10], Heisenberg MPO [89] and PEPO time-evolution [90], gPEPS [91], as well as noisy [92] and sparse Pauli dynamics [93] methods can solve the same problem more efficiently, and can even produce more accurate results than the quantum processor itself.
Another important class of experiments suitable for TN techniques is a random circuit sampling (RCS). RCS is a computational task where a quantum computer executes a random $n$-qubit quantum circuit and output the measurement results (a length-$n$ bitstring) of the resulting quantum state. This task is classically hard because the existence of an efficient classical algorithm for estimating the probability of a given output would lead to the collapse of the polynomial hierarchy [94, 95]. As a result, RCS has been experimentally demonstrated many times as a proof of quantum computational advantage [96, 97, 98, 57, 18]. Additionally, applications based on RCS in which classical simulation using TNs consistitute an integral part have been proposed [99, 100].
The leading approach for simulating RCS with the lowest computational cost and the highest fidelity is contracting the TN representing the quantum circuit as a network of contracted tensors that correspond to the applied gates [17]. Contracting a fraction of slices effectively performs a finite fidelity simulation [56]. Approximation and simplification of the original TN could also be implemented alongside slicing, which reduces the fidelity [67]. Additionally, conventional RCS experiments usually request millions of samples from a single circuit. One can reduce the cost by leaving some indices open [101, 102, 7], using a sparse output state [67], or reuse intermediate results [56, 66, 7]. One can also perform some post processing on the samples to spoof the quality metric [69].
Another commonly analyzed approach is the MPS based on DMRG [103]. This approach also breaks the TN corresponding to a sample amplitude into three parts: the beginning of the circuit with fixed input indices, the middle, and the end with fixed output indices. The first and third parts can be represented using an MPS, and the amplitude can be obtained by contracting the two MPSs with the middle part in between.
Finally, prior to the development of the aforementioned methods, other TN techniques such as MPO [104], MPDO [105], PEPS [106], conversion of 3D to 2D networks [63] are utilized.
@@ III.2 Analog evolution
Analog quantum computers are devices where a specific class of models is implemented “natively” on the quantum hardware. These systems include optical lattices of neutral toms [107], trapped ions [108], Rydberg tweezers [109], photonic waveguides [110], as well as superconducting circuits [111]. Contrary to digital quantum processors, analog quantum devices allow the simultaneous, time-dependent continuous control of pairwise interactions of all the qubits available on the quantum chip. The major drawbacks of analog quantum simulators are the calibration errors of Hamiltonian parameters and decoherence. One also needs to make sure that the quantum processor is faithfully behaving according to closed-system Schrödinger time-evolution for the largest number of qubits possible. TNs provide an invaluable benchmarking tool in this regard.
@@
A recent simulation technique was able to achieve comparable accuracies to the quantum annealer [115] for two- and three-dimensional systems. It uses belief propagation for time evolution, and more sophisticated variants of belief propagation for calculating expectation values. However, this method is not immediately appicable to the inifite dimensional biclique lattice studied by the quantum annealer.
@@ III.3 Boson sampling
Since the transmission matrices of boson sampling experiments are approximately Haar-random [117, 118, 119, 120], using an MPS approach is suitable despite the high dimensionality of the optical interferometer [121, 122]. Additionally, for simulations without photon loss, photon number conservation implies a $U(1)$ symmetry, which is further exploited to reduce the cost of the MPS [121, 122]. In reality, various experimental noises, especially photon loss, scale with the system size, and the complexity-theoretic argument is not applicable in this case. Therefore, directly approximating the mixed state in the Fock basis using an MPO can be efficient. In this approach, if we have $N$ input photons and $N_{\rm out}\propto N^{\gamma}$ output photons (where $\gamma$ is the scaling exponent), the MPO entanglement entropy (MPO EE) grows as $S=O(N^{2\gamma-1}\log N)$, which indicates a logarithmic growth of the MPO EE when $N_{\rm out}\propto\sqrt{N}$[122, 123].
Further, for Guassian boson sampling [124], the lossy output states can be modeled by applying random classical operations on a pure state [9, 125]. This allows the simulation of the pure state with an MPS. Further, one can optimize this pure state to have significantly fewer photons compared to the final output state [9]. The bond dimension grows logarithmically when $N_{\rm out}\propto\sqrt{N}$, and this theoretically guarantees polynomial growth of the bond dimension to fixed fidelity, allowing the largest boson sampling experiments to be simulated on up to $288$ GPUs in under two hours [9].
Additionally, one can simulate boson sampling in the Heisenberg picture in a way similar to the simulation of quantum circuits. Instead of evolving the quantum state using a TN, this approach evolves the Fock state projector of the desired measurement outcome [126]. For Gaussian boson sampling, the Schrödinger picture approach evolves a Gaussian state, which usually has a larger bond dimension, whereas the Heisenberg picture evolves a Fock state.

### v2

# thanks: Equal contributionthanks: Equal contribution

@@ Abstract
In the rapidly evolving field of quantum computing, tensor networks serve as an important tool due to their multifaceted utility. In this paper, we review the diverse applications of tensor networks and show that they are an important instrument for quantum computing. Specifically, we summarize the application of tensor networks in various domains of quantum computing, including simulation of quantum computation, quantum circuit synthesis, quantum error correction and mitigation, and quantum machine learning. Finally, we provide an outlook on the opportunities and the challenges of the tensor-network techniques.
@@ I Introduction
Simulation of quantum computation (Section III) Unstructured TNs, matrix product states (MPS), projected entangled pair states (PEPS), tree tensor network states (TTNS), multi-scale entanglement renormalization ansatz (MERA), density matrix renormalization group (DMRG), various renormalization methods Simulation of quantum circuits, analog quantum processors, boson sampling, quantum algorithm benchmarking, quantum many-body physics simulation Simulation complexity reduction, efficiency in representation of some quantum states Representing highly-entangled quantum states, simulating long-time dynamics, simulating deep quantum circuits
@@
Quantum error correction (QEC) and mitigation (QEM) (Section V) Tensor-network decoders, TN inspired quantum error-correcting codes Tensor-network decoders of error-correcting codes Efficiency and interpretability in decoding, interpretability in creating and studying error-correcting codes, reduction in error mitigation overhead Computational resources overhead, speed, lack of open-source software
Quantum machine learning (QML) (Section VI) Variational tensor-network circuits as quantum neural networks (QNNs) Generative and discriminative models, quantum data classification, quantum data encoding Analytical interpretability, qubit efficiency, noise resilience, absence of barren plateaus, potential reduction in parameters number, avoiding costly TN contraction on classical computer Requirement for nonlocal interactions in quantum hardware for many TN circuits, lack of theoretical guarantees for advantage
@@
Tensor networks (TNs) have become a useful tool in many areas of physical and mathematical sciences, especially in the field of quantum information science. The interest in quantum computing (QC) has driven a lot of the development in TNs because they are used to represent and manipulate quantum states and processes.
@@
Additionally, TNs provide a general framework of analyzing mathematical objects frequently encountered in quantum information science, which makes them attractive tools to address many other challenges faced by QC. Table 1 provides an overview of TN applications in QC that we will discuss in this paper and spans four subdomains: simulation of quantum computation, quantum circuit synthesis (QCS), quantum error correction (QEC) and mitigation (QEM), and quantum machine learning (QML). The Table 1 outlines the specific methods used, their primary applications, key advantages, and the challenges they encounter.
We structure the rest of the paper as follows. Section II briefly introduces tensor networks, key building blocks, and commonly used methods. Sections III, IV, V and VI discuss the aforementioned domains of application. Finally, we conclude the paper with discussions on the overall advantages of tensor networks, and provide an outlook of the future on how tensor networks may benefit quantum computing in Section VII.
@@ II Tensor-network methods
where $\Sigma$ is a diagonal matrix of the singular values and could be absorbed into $U$ or $V$, and $U,V$ are isometries ($U^{\dagger}U=I,V^{\dagger}V=I$). This decomposition presents an opportunity for approximate representation of the original tensor by trimming the singular values, either by keeping only the $k$ largest singular values or discarding singular values smaller than some threshold. Furthermore, the Eckart–Young theorem [9] states that for approximations with a fixed rank, the solution provided by SVD is optimal. There exist other matrix decompositions which we do not discuss for the sake of brevity [10].
Graphically, the SVD decomposition is shown in Figure 1(e). There, isometries are represented as orangle nodes with directional bonds shown in Figure 1(f). For isometries $U$, although $U^{\dagger}U=I$, but $UU^{\dagger}\neq I$ unless $U$ is also unitary. Therefore, the arrow notation is used to differentiate the indices of the isometry. Unitaries, being a special case of isometries, are represented as magenta nodes with directional bonds as shown in Figure 2.
II.1 Common ansatzes
Figure 2: Different types of TNs: (a) the matrix product state, (b, c) the projected entangled pair states in 2D and 3D, (d) the tree tensor network state, and (d) the multiscale entanglement renormalization ansatz. The orange circles represent isometric tensors and the magenta circles represent unitary disentanglers.
The matrix product state (MPS) is the simplest yet the most widely used ansatz (in this context, this term refers to the structure as well as constraints on the TN’s tensors). In general, the amplitudes of an $N$-body quantum wavefunction form an $N$th-order tensor, the MPS provides an approximate representation of this tensor as
@@ II.1 Common ansatzes
which is a contraction of the $\Gamma_{i}$ tensors, each corresponding to a site in the system as shown in Fig. 2 (a). Here, $\chi$ is the parameter called the bond dimension which controls the accuracy of compression. Closely related to MPS is the matrix product operator (MPO) concept, which uses a very similar representation for many-body operators instead of vectors (which is the case for the MPS). This is accomplished by adding an additional dual index to each $\Gamma_{i}$ tensor.
This TN is especially effective for simulating one-dimensional (1D) quantum systems. It follows the area law of entanglement [11] (the entanglement between two parts of a many-body systems growing at most as the size of the boundary), but has been used for other systems as well [12, 7]. Its contraction cost is polynomial in the system size, meaning the physical observables can be evaluated efficiently. This is because local physical observables can be expressed as a tensor network and can be evaluated by contracting it. The concept of MPS has independently emerged in computational mathematics, where it is called the tensor train [13].
Projected entangled pair states (PEPS), shown in Fig. 2 (b) and Fig. 2 (c), aim at generalizing the 1D locality of the MPS to higher dimensions, making them capable of reproducing large entanglement and suitable for simulating high dimensional quantum systems. However, algorithms for PEPS are typically much more computationally demanding [3, 14].
Tree Tensor Networks States (TTNS)[15, 16], shown in Fig. 2 (d) generalize the MPS from local to nonlocal correlations. MPS can be understood as the special case of a TTNS with a maximally unbalanced tree topology. Balanced tree topologies provide a worst-case correlation length that is logarithmic in the number of leaves and thereby allow to capture long-range interactions. The scaling of TTNS is comparable to that of MPS[16], however, the rank required for accuratly describing the state might scale rapidly. TTNS have been independently developed in the chemical physics community under the name multilayer multiconfigurational Hartree[17].
The multiscale entanglement renormalizaton ansatz (MERA), shown in Fig. 2 (e), is designed to capture the entanglement structure of quantum systems across different length scales and can reproduce logarithmic violation of the area law.
@@ II.2 Manipulation of TNs
While TNs can be used to represent a broad range of mathematical objects in quantum information science, their popularity stems primarily from their power as a tool for obtaining meaningful, interpretable quantities such as operator expectation values, reduced density matrices, amplitudes, samples, and more. Due to the high interest in the TN techniques, a plethora of software libraries suitable for different types of TN manipulations are available, such as cuQuantum [20], quimb [21], ITensor [22], and many more [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@ II.2.1 Update
In addition to the DMRG algorithm, gradient-based optimization methods can also be employed if the gradient of the objective function (e.g., energy) is accessible, for instance, through automatic differentiation [40, 41]. Additionally, Riemannian optimization [42, 43] can be particularly helpful when one needs to maintain the isometric properties of the TN [44, 45].
@@
One may also consider the time evolution of a quantum state or an operator in the TN representation. For discrete time steps, this can be done by applying local gates or Kraus operators through contraction, and restore the original form of the TN with SVD. These approaches are called time-evolving block decimation (TEBD). Continuous time evolution under a Hamiltonian can be approximated by discrete steps and the discrete method can be applied if the Hamiltonian is sufficiently short ranged. For long-range Hamiltonians, TEBD becomes inefficient, and methods such as MPO $W^{II}$, TDVP, and Krylov approaches are preferred; see [48] for a recent review. Krylov methods use Lanczos diagonalization (an approximate method) to compute the action of the time evolution operator, while TDVP evolves the state directly on the MPS manifold for fixed bond dimension [49, 50, 51, 52].
@@ II.2.2 Contraction
Finding the optimal contraction scheme is itself an NP-complete problem [54, 55]. However, much progress has recently been made using heuristic approaches targeting the total cost of the contraction operation. Specifically, these methods have been based on recursive graph partitioning [56, 23, 57], simulated annealing [58, 59] and now reinforcement learning [60, 61]. The extraordinary sensitivity of the cost to the contraction tree quality has led to improvements by many orders of magnitude for some problems such as, for example, quantum circuit simulation [62, 23, 63].
For exact simulation of quantum circuits, two additional techniques are central to achieving a state-of-the-art performance. First, whilst the time cost of a contraction might be acceptable, the space cost in terms of intermediate memory might still be enormous. Slicing (also known as cutting or projecting) splits a contraction into many smaller independent contractions [64, 65, 66, 67, 68], each of which for example can fit on a GPU and be performed in a massively parallel way [69, 58]. Depending on the geometry, this can sometimes be performed with very little overhead. Secondly, if one wants to evaluate many related tensor networks differing only in some entries (for example, a set of amplitudes of different output basis states of a single quantum circuit), one can employ multi-contraction [62, 5] to cleverly avoid repeating the same computation. The largest exact simulations of quantum circuits use both of these techniques in tandem [5, 70, 71, 72].
@@ III.1 Gate-based quantum computation
A good example is the simulation of quantum circuits used to simulate many-body physics, such as the experiment on the dynamics of a kicked Ising model [90]. While the problem was considered initially intractable using state-vector MPS and isometric-TN [91] approaches, it was later demonstrated that many other TN methods [92, 93, 41, 94, 95] can solve the same problem more efficiently, and can even produce more accurate results than the quantum processor itself.
Another important class of experiments suitable for TN techniques is random circuit sampling (RCS). RCS is a computational task where a quantum computer executes a random $n$-qubit quantum circuit and output the measurement results (a length-$n$ bitstring) of the resulting quantum state. This task is classically hard because the existence of an efficient classical algorithm for estimating the probability of a given output would lead to the collapse of the polynomial hierarchy [96, 97], which theoretical computer scientists believe to be highly unlikely. As a result, RCS has been experimentally demonstrated many times as a proof of quantum computational advantage [98, 99, 100, 59, 19]. Additionally, applications based on RCS in which classical simulation using TNs constitute an integral part have been proposed [101, 102, 103, 104].
The leading approach for simulating RCS with the lowest computational cost and the highest fidelity is contracting the TN representing the quantum circuit as a network of contracted tensors that correspond to the applied gates [18]. Contracting a fraction of slices effectively performs a finite fidelity simulation [58]. Approximation and simplification of the original TN could also be implemented alongside slicing, which reduces the fidelity [70].
Additionally, conventional RCS experiments usually request millions of samples from a single circuit. One can reduce the cost by leaving some indices open [105, 106, 5], using a sparse output state [70], or reuse intermediate results [58, 62, 5]. One can also perform some post processing on the samples to spoof the quality metric [72].
Another commonly analyzed approach is the MPS based on DMRG [107]. This approach also breaks the TN corresponding to a sample amplitude into three parts: the beginning of the circuit with fixed input indices, the middle, and the end with fixed output indices. The first and third parts can be represented using an MPS, and the amplitude can be obtained by contracting the two MPSs with the middle part in between. Noteworthy, this method has been recently used to prove quantum utility in simulating the quantum Ising model’s dynamics [108].
Finally, prior to the development of the aforementioned methods, other TN techniques such as MPO [109], MPDO [110], PEPS [111], conversion of 3D to 2D networks [67] and TTNS [112, 113] were utilized to simulate gate-based quantum computations. Overall, while the earlier RCS experiments [98, 99, 100] can now be simulated, the more recent experiments with better fidelity and larger circuit volume still remain hard [59, 19, 114].
@@ III.2 Analog evolution
Analog quantum computers are devices where a specific class of models is implemented “natively” on the quantum hardware. These systems include optical lattices of neutral atoms [115], trapped ions [116], Rydberg tweezers [117], photonic waveguides [118], as well as superconducting circuits [119]. Contrary to digital quantum processors, analog quantum devices allow the simultaneous, time-dependent continuous control of pairwise interactions of all the qubits available on the quantum chip. The major drawbacks of analog quantum simulators are the calibration errors of Hamiltonian parameters and decoherence. One also needs to make sure that the quantum processor is faithfully behaving according to closed-system Schrödinger time-evolution for the largest number of qubits possible. TNs provide an invaluable benchmarking tool in this regard.
@@
A recent simulation technique was able to achieve comparable accuracies to the quantum annealer [123] for two- and three-dimensional systems. It used belief propagation for time evolution, and more sophisticated variants of belief propagation for calculating expectation values. However, no such simulation was performed for the infinite dimensional biclique lattice studied by the quantum annealer [124].
@@ III.3 Boson sampling
Since the transmission matrices (matrices describing the relationship between input and output optical modes) of boson sampling experiments are approximately Haar-random [126, 127, 128, 129], using an MPS approach is despite the high dimensionality of the optical interferometer [130, 131]. Additionally, for simulations without photon loss, photon number conservation can be exploited to reduce the cost of the MPS [130, 131].
In reality, various experimental noises, especially photon loss, scale with the system size, and the complexity-theoretic argument is not applicable in this case. Therefore, directly approximating the mixed state in the Fock basis using an MPO can be efficient. In this approach, if we have $N$ input photons and $N_{\rm out}\propto N^{\gamma}$ output photons (where $\gamma$ is the scaling exponent), the MPO entanglement entropy (MPO EE, which roughly characterizes the simulation hardness) grows as $S=O(N^{2\gamma-1}\log N)$, which indicates a logarithmic growth of the MPO EE when $N_{\rm out}\propto\sqrt{N}$[131, 132].
Further, for Guassian boson sampling [133], the lossy output states can be modeled by applying random classical operations on a pure state [7, 134], allowing the state to be simulated with an MPS. Further, one can optimize this pure state to have significantly fewer photons compared to the original state [7]. The bond dimension grows logarithmically when $N_{\rm out}\propto\sqrt{N}$, and this theoretically guarantees polynomial growth of the bond dimension to fixed fidelity. This enabled the largest boson sampling experiments to be simulated on up to $288$ GPUs in under two hours [7], meaning that no boson sampling experiments demonstrate clear evidence for beyond classical hardness.
Additionally, one can simulate boson sampling in the Heisenberg picture in a way similar to the simulation of quantum circuits. Instead of evolving the quantum state using a TN, this approach evolves the Fock state projector of the desired measurement outcome [135]. For Gaussian boson sampling, the Schrödinger picture approach evolves a Gaussian state (an infinite superposition of Fock states), which usually has a larger bond dimension, whereas the Heisenberg picture evolves a Fock state.